OpenStudy (anonymous):

Solve for x. x-7/8 = 7/2 A. x=35 B. X=9(2/7) C.x= 31(1/2) D. x=63

OpenStudy (anonymous):

\[x-\frac{ 7 }{ 8 }=\frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

First,add 7/8 on both sides.

OpenStudy (anonymous):

\[x-\frac{ 7 }{ 8 }+\frac{ 7 }{ 8 }=\frac{ 7 }{ 2 }+\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

\[x=\frac{ 7 }{ 2 }+\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Then make the denominator the same

OpenStudy (anonymous):

14/ 10?

OpenStudy (anonymous):

\[x=\frac{ 7\times4 }{ 2\times4 }+\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Okay then what do I do next?

OpenStudy (anonymous):

\[x=\frac{ 28 }{ 8 }+\frac{ 7 }{ 8 }\]\[x=\frac{ 28+7 }{ 8 }\]Add up

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

\[x=\frac{ 35 }{ 8 }\]\[x=?\]make it into proper fraction

OpenStudy (anonymous):

The options are wrong

OpenStudy (anonymous):

It's not wrong

OpenStudy (anonymous):

Pls recheck ur question...maybe u written it wrongly

OpenStudy (anonymous):

They are fractions

OpenStudy (anonymous):

\[\frac{ x-7 }{ 8 }=\frac{ 7 }{ 2 }\]like this?

OpenStudy (anonymous):

First,multiply both sides by 8. Then,add 7 on both sides.

OpenStudy (anonymous):

\[(\frac{ x-7 }{ 8 })\times8=(\frac{ 7 }{ 2 })\times8\]\[x-7=28\]\[x-7+7=28+7\]\[x=?\]solve it @jojoheart